OpenStudy (anonymous):

Find the area of the regular polygon. Give the answer to the nearest tenth. dodecagon with a perimeter of 108 cm

OpenStudy (mathmate):

Area of a polygon = (perimeter/2)*distance from centre to one of the sides. A dodecagon has 12 sides. Each side = perimeter/12 = 9 The exterior angle is 360/12=30 degrees interior angle is 180-30=150 deg. Base angle of isosceles triangle formed with the base and the centre = 150/2=75 degrees Height of triangle = (9/2) tan(75 deg.) Area of dodecagon =(108/2)*4.5tan(75deg.) = 907 approx.